OpenStudy (anonymous):

Given f (x) = 7x^2 + 4x - 13, A = sum of roots of f (x) = 0, B = product of roots of f (x) = 0, find ( A + B)^2

OpenStudy (anonymous):

Won't factor, use quadratic formula

OpenStudy (anonymous):

not to butt in but you can find the sum of the roots and the produce of the roots without finding the roots

OpenStudy (anonymous):

Alright then, the floor is yours lol

OpenStudy (anonymous):

you can also prove those formulas using the quadratic formula, but sum of roots of a quadratic \[ax^2+bx+c=0\] \[r_1+r_2=-\frac{b}{a}\] produce is \[r_1r_2=\frac{c}{a}\]

OpenStudy (anonymous):

*product

OpenStudy (anonymous):

The answer is 289/49. Could you please show me the steps?

OpenStudy (anonymous):

product of the roots, according to satellite's formula is -13/7

OpenStudy (anonymous):

sum is -4/7

OpenStudy (anonymous):

\[(-\frac{13}{7}-\frac{4}{7})^2=(-\frac{17}{7})^2=\frac{289}{49}\]

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

could you explain how did you get the A .